# Parallel Batch Scheduling With Incompatible Job Families Via Constraint Programming

### v2

@@ Abstract
This paper addresses the incompatible case of parallel batch scheduling, where compatible jobs belong to the same family, and jobs from different families cannot be processed together in the same batch. Existing constraint programming (CP) models for this problem fail to synchronize the processing of the jobs within their batch, resulting in batch interruptions. In the context of the diffusion area in the semiconductor manufacturing process, these interrupted solutions would disrupt the thermal stability required for a uniform dopant distribution on the wafers. This paper proposes three new CP models that directly tackle these interruptions in the formulation, including two adaptions of existing models and a novel Redundant Synchronized (RS) model. These existing and novel models are compared on standard test cases, demonstrating the superiority of the RS model in finding optimal or near-optimal solutions quickly.
@@ I Introduction
A more complex p-batch problem arises in the front-end of the SM process, in the diffusion area of the wafer fab, where dopants are introduced into the silicon wafers at high temperatures to alter their electrical properties. This is achieved using diffusion furnaces that operate in batches, applying constant heat on the wafers for long periods, typically 10-12 hours [17]. These prolonged processing times are essential to ensure the thermal stability required for an even distribution of dopants on the wafers [18]. Due to the chemical nature of the process, not all the jobs can be batched together. Instead, incompatible job families are defined by the distinct “recipes” of the jobs, which dictate unique dopants and thermal conditions [2]. Jobs with the same recipe are said to be compatible and therefore grouped into the same family; jobs in the same family also have the same processing time [1, 19]. Jobs of different families are said to be incompatible and cannot be processed together in the same batch, as this may result in unwanted cross-contamination in the chemical process. Hence, the scheduling problem in the diffusion area belongs to the incompatible case of p-batch, where the batches can only contain jobs of the same family [2, 18]. Furthermore, due to the risk of burning the wafers or damaging the chemical process, jobs cannot remain in the furnace beyond their processing time. Moreover, the processing of a batch cannot be interrupted to include or remove jobs mid-process, as this would disrupt the thermal stability and jeopardize the uniformity of the diffusion [18]. Thus, jobs inside a batch must be synchronized, i.e., their processing must start and end together. This synchronization of jobs within their batch is critical for the diffusion but is typically neglected in traditional p-batch models, like the ones for the compatible case of p-batch in the burn-in operations.
Existing exact methods in the literature for p-batching include mixed-integer programming (MIP) [6, 18] and Constraint Programming (CP) models [8, 18]. MIP models do not scale well and are outperformed by the CP approaches [2, 18]. However, the existing CP models fail to synchronize the jobs within their batch, allowing interruptions mid-process. This paper remedies this shortcoming by proposing three new CP models that directly tackle these interruptions in their formulation, and explores the impact of breaking symmetries in the solution. Furthermore, the paper conducts thorough computational experiments that demonstrate the benefits of the proposed models over those in the literature.
@@
The remainder of this paper is organized as follows. Section II presents a literature review on p-batching. Section III clearly outlines our contributions. Section IV formally presents the description of the problem we address. Section V presents an existing CP model that allows interruptions. Section VI presents our proposed CP models that remedy them. Section VII presents our computational experiments and their results. Finally, Section VIII presents the conclusions and outlines future lines of research.
@@ II Literature Review
The standard classification scheme for scheduling problems provided by Graham et al. [20] is $\alpha|\beta|\gamma$, where $\alpha$ denotes the scheduling environment, $\beta$ describes the job and family characteristics and any restrictions, and $\gamma$ defines the objective function. This section focuses on relevant articles that deal with multiple parallel machines (i.e., $\alpha=P$). For a more comprehensive overview of single-machine p-batch, interested readers are referred to the recent survey by Fowler and Mönch [2]. For $\beta$, the focus is on articles that deal with multiple incompatible families (i.e., $P|p-batch,incompatible|\gamma$), although key articles with a single compatible family (i.e., $P|p-batch|\gamma$) are also considered, as the paper extends some of their concepts to the incompatible case. The review also includes some articles that consider non-identical job sizes ($s_{j}$) and release times ($r_{j}$). For $\gamma$, the reviewed objective functions primarily fall into three categories: (i) utilization objectives such as minimizing the makespan; (ii) cycle time objectives such as minimizing the total completion time (TCT) and its weighted counterpart (TWCT); and (iii) on-time delivery objectives such as minimizing earliness, lateness, tardiness, number of tardy jobs, including their weighted variants [2]. However, since the diffusion area is usually a bottleneck in the wafer fab, fab leaders often focus on the cycle time of the jobs to increase the effective movement of products [19]. Thus, the focus is on minimizing the TWCT, which specifically targets the cycle time performance [2].
@@
Most of the articles that consider multiple parallel machines usually focus on a single family (i.e., $P|p-batch|\gamma$) [2]. This problem is NP-hard [21, 2] and multiple heuristic algorithms [22, 23, 24] as well as meta-heuristics [25, 26, 27] have been proposed. Some exact approaches include mixed-integer linear programming (MILP) models [16] and efficient arc-flow formulations [28, 29, 30, 5]. Recently, CP has been used for this problem in contexts different from SM, such as the composite manufacturing industry [11], and AM processes where parts with irregular shapes are 3D printed [8, 9, 10]. Nascimento et al. [8, 9] presented an Assign-and-Schedule (AS) CP model that uses binary variables for the assignment of the parts to printing batches and CP scheduling logic to sequence the batches on the 3D printers. Their assignment logic already locates the parts inside the batch, so they do not explicitly include capacity constraints. However, they only consider compatible jobs (i.e., one family). This paper extends their formulation to handle multiple incompatible families, and include additional constraints to handle the batch size requirements.
@@
The literature that deals with multiple parallel machines and incompatible job families (i.e., $P|p-batch,incompatible|\gamma$) is rather scarce [2]. Since this problem is NP-hard [4], the literature is dominated by heuristics and meta-heuristics [4, 31, 32, 33, 34, 35, 19]. However, exact methods have also been proposed, such as DP [19] and MIPs [6, 18]. Cakici et al. [6] proposed a time-indexed (TI) MIP model inspired by the truck delivery problem. They also presented two variable-neighborhood search (VNS) heuristics that outperform the MIP model. Later, Ham et al. [18] presented another MIP formulation based on positional assignment (PA) variables in the context of the diffusion area. They also proposed a CP model that outperforms both the PA and the TI MIP models, as well as the VNS heuristics. However, this CP model fails to synchronize the jobs within their batch, allowing job interruptions mid-process. This paper addresses this issue and proposes a set of synchronized CP models that tackle the interruptions directly in the formulation.
@@ III Contributions
As evidenced in the previous section, most of the existing CP models for p-batch only deal with a single compatible family [8, 9, 10]. Among these, the only model that synchronizes the jobs relies on a constraint specifically designed to ensure the correct batch duration in the AM context, which does not directly translate to the diffusion problem [10]. Moreover, the existing CP model that considers multiple incompatible families fails to synchronize the jobs within their batch, allowing interruptions mid-process [18].
This paper addresses these gaps in the literature by proposing three new CP models that consider multiple incompatible job families and directly tackle the interruptions in their formulation while synchronizing the jobs within their batch: (i) An Assign-and-Schedule (AS) model inspired by Nascimento et al. [8, 9] that handles the assignment of jobs to batches with binary variables, and the sequencing of batches on machines using CP scheduling logic; (ii) A Synchronized (S) model inspired by Cakici et al. [10] that replaces the binary assignment with CP scheduling logic, and guarantees the presence of the required batches with an implication constraint; (iii) and a new Redundant Synchronized (RS) model that includes redundant variables and constraints that help the search process and improves computational and quality performance. Furthermore, the paper explores the impact of symmetry-breaking constraints in the search process. The computational experiments demonstrate the superiority of the RS model over existing models.
@@ IV Problem Description
Let $\mathcal{J}$ be the set of jobs, $\mathcal{F}$ be the set of incompatible job families, and $\mathcal{M}$ be the set of machines. All jobs can be scheduled on all machines. Compatible jobs belong to the same family. Hence, let $f_{j}\in\mathcal{F}$ be the family of job $j\in\mathcal{J}$, and $\mathcal{J}_{f}=\set{j\in\mathcal{J}:f_{j}=f}$ be the subset of jobs that belong to family $f\in\mathcal{F}$. All the jobs of the same family have the same processing time, thence, let $p^{f}$ be the processing time of family $f$ and let $p_{j}=p^{f}$ be the processing time of job $j\in\mathcal{J}_{f}$. Additionally, each job $j$ has a size $s_{j}$ (i.e., number of parts), a weight $\omega_{j}$ (i.e., priority), and a release time $r_{j}$. Each machine $m\in\mathcal{M}$ can simultaneously process multiple jobs of the same family $f\in\mathcal{F}$ in batches whose size must be between $l_{f}$ and $u_{f}$, as evidenced in the instances of the SM testbed 2020 [17]. Jobs in the same batch have to start and end together in a synchronized fashion, and jobs of different families cannot be processed together in the same batch. The goal is to minimize the total weighted completion time (TWCT) of the jobs. In the classification by Graham et al. [20], this problem is $P|s_{j},r_{j},p-batch,incompatible,synchronized|TWCT$.
V The Existing CP Model
CP is a powerful method for solving combinatorial problems by defining constraints that a solution must satisfy. In CP, variables are assigned values from their domains, and the system works to ensure that all constraints are met. The search process explores the solution space in the form of a tree, where fixing the values of variables at each node branches out potential solutions. A key aspect of CP is the use of constraint propagation, which reduces the search space by pruning values from variable domains that cannot satisfy the constraints. Backtracking occurs when a branch leads to an invalid solution, allowing the search to revert to an earlier state and explore different variable assignments [36].
Scheduling is one of the most successful application areas of CP. It leverages interval variables to represent tasks or operations over time. These interval variables encapsulate three key components: the start time, the end time, and the presence status of the interval (indicating whether the task is executed). The start time and the end time define the size of the interval. The presence attribute allows CP to model both mandatory and optional activities within a schedule. In particular, it plays a crucial role in handling optional tasks and alternative resource allocations. Interval sequence variables allow to model sequences of tasks. CP also uses cumulative functions to model the evolution of resource usage over time, ensuring that resource capacities are respected by aggregating the resource demands of all overlapping tasks. These cumulative functions are crucial in preventing over-utilization of resources like machines, personnel, or energy. The efficiency of the search process is further enhanced by global constraints that exploit the structure of the problem to propagate information and prune variable domains efficiently. Together, CP and its tools for constraint propagation, domain pruning, backtracking, and cumulative functions provide a structured and efficient approach for solving complex scheduling problems [36, 37]. The interested reader on an in-depth CP overview is referred to the well-known book by Van Beek et al. [36], and the paper by Laborie et al. [37] about IBM ILOG CP Optimizer.
@@ V The Existing CP Model
Model 1 presents the CP model proposed by Ham et al. [18] to schedule incompatible jobs in the diffusion area. Equation (1) defines an interval variable $x_{j}$ of the form $[a,a+p_{j})$ (i.e., it has a duration of exactly $p_{j}$ units of time), which represents job $j\in\mathcal{J}$. Note that this interval can only start after the job’s release time. Equation (1) defines an optional interval variable $x_{jm}$, also of size $p_{j}$, that represents the option of processing job $j$ on machine $m\in\mathcal{M}$. The presence of the interval $x_{jm}$ is optional since it is allowed to take the value $\perp$, which indicates its absence in the solution. These optional interval variables allow the model to decide which machine should process the jobs. Equation (1) defines a function ${\mathfrak{f}}_{m}$, which is the unique state of the machine $m$ at every point in time. Equation (1) defines a cumulative function ${\mathscr{l}}_{mf}$, which represents the load on machine $m$ of jobs of the family $f\in\mathcal{F}$. It uses the pulse function, which receives two inputs: An interval $x_{jm}$ and an integer value $s_{j}$. This pulse function tells the cumulative function to increment its value in $s_{j}$ units during the interval $x_{jm}$, if present. Hence, the function ${\mathscr{l}}_{mf}$ accumulates the sizes of overlapping jobs of family $f\in\mathcal{F}$ on machine $m$ at every point in time.
Model 1 Non-synchronized model by Ham et al. [18]
@@
$x_{j}$$\in\{[a,a+p_{j}):a\in\mathbb{Z},r_{j}\leq a\},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J};$ $(1)$
$x_{jm}$$\in\{[a,a+p_{j}):a\in\mathbb{Z}\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M};$ $(2)$
@@
${\mathscr{l}}_{mf}$ $=\sum_{j\in\mathcal{J}_{f}}\text{pulse}\left(x_{jm},s_{j}\right),$ $\forall\leavevmode\nobreak\ m\in\mathcal{M},f\in\mathcal{F}.$ $(4)$
@@
alwaysEqual $\left({\mathfrak{f}}_{m},x_{jm},f_{j}\right);$
$\forall\leavevmode\nobreak\ m\in\mathcal{M},f\in\mathcal{F},j\in\mathcal{J}_{f}:$ (8)
alwaysIn $\left({\mathscr{l}}_{mf},x_{jm},l_{f},u_{f}\right).$
Objective function (1) minimizes the TWCT. Constraints (1) use the alternative global constraint, which receives two inputs: an interval variable $v=x_{j}$, and a set of optional interval variables $B=\set{x_{jm}}_{m\in\mathcal{M}}$. This global constraint ensures that, if the interval variable $v$ is present, then only one of the optional interval variables in $B$ gets selected to be present, and $v$ gets synchronized with the selected interval in $B$ [38]. Hence, constraints (1) enforce each job to be processed on exactly one machine. Constraints (1) use the global constraint alwaysEqual, which receives three inputs: A state function $h={\mathfrak{f}}_{m}$, an interval variable $v=x_{jm}$, and an integer value $a$. This global constraint ensures that the state function $h=a$ during the interval $v$, if present [38]. Hence, constraints (1) ensure that jobs of only one family are processed at any time on the machines. Finally, constraints (1) use the global constraint alwaysIn, which receives four inputs: A cumulative function $g={\mathscr{l}}_{mf}$, an interval variable $v=x_{jm}$, and integer values $a=l_{f}$ and $b=u_{f}$. This global constraint ensures that $a\leq g\leq b$ during the interval $v$, if present [38]. Hence, constraints (1) guarantee that the load on the machines is within the required batch size limits of the family being processed.
V-A Shortcomings of the Existing CP Model
Model 1 clearly selects one machine for each job, guarantees that only jobs of one family are processed at the same time on each machine, and ensures that the load on each machine is within the batch size limits of the family being processed on it. Nonetheless, the model never ensures that the jobs being processed on the machines at any point in time start and end together in a synchronized fashion. Model 1 is called Non-synchronized (NS), as it allows job interruptions mid-process.
Consider the scenario in Table I as an illustrative example with four jobs, one family, and one machine. Assume that the size of each batch must be between $l_{1}=50$ and $u_{1}=100$. Figure 1 shows the Gantt chart of the optimal solution achieved by the NS model. The thick blue dotted line shows the load on the machine with jobs of family $1$. This solution has a TWCT $=1,550$. The state of the machine is the family of the jobs being processed, representing a batch of jobs of family $1$. However, the jobs are not synchronized, allowing two batch interruptions and four job interruptions. At time 12, jobs 1 and 3 are interrupted to include jobs 2 and 4. At time 15, jobs 2 and 4 are interrupted to remove jobs 1 and 3. If these interruptions happen in the diffusion area, the thermal stability required by the wafers would be damaged.
@@ V-A Shortcomings of the Existing CP Model
Job $j$ Size $s_{j}$ Weight $\omega_{j}$ Release time $r_{j}$ Family $f_{j}$ Processing time $p_{j}$
@@
Figure 1: Optimal solution of the Non-synchronized model
V-B Feasibility Restoration
The NS model allows job interruptions, making the solutions infeasible to the problem. To remedy this, after retrieving the solution of the NS model, this paper proposes to conduct a Feasibility Restoration (FR) step as follows: the scheduled jobs on each machine are sorted by their start time, and the FR iterates through them while keeping track which jobs start, which jobs are currently in process, and which jobs are completed. If there are jobs currently in process when a new job starts, FR either pulls the interrupted jobs to start with the new ones, or pushes the starting jobs right after the jobs that are currently in process (and shifts all the future jobs accordingly to maintain the sequence). This local decision of pulling or pushing jobs is made greedily while ensuring feasibility of the batch size limits and compatibility of batched jobs. By the end of the process, the solution is guaranteed to be feasible without interruptions. Nonetheless, no guarantee on the quality of the resulting solution can be made. To remedy this, Section VI proposes three new CP models that directly tackle interruptions in their formulations.

### v3

@@ Abstract
This paper addresses the incompatible case of parallel batch scheduling, where compatible jobs belong to the same family, and jobs from different families cannot be processed together in the same batch. The state-of-the-art constraint programming (CP) model for this problem relies on specific functions and global constraints only available in a well established commercial CP solver. This paper expands the literature around this problem by proposing four new CP models that can be implemented in commercial and open-source solvers: a new model that relies on automaton constraints, and three alternative models that integrate assignment and scheduling decisions with different strategies and global constraints. Extensive computational experiments on standard test cases under multiple objectives and multiple solvers demonstrate the implementation flexibility and competitive performance of the proposed models.
@@ I Introduction
A more complex p-batch problem arises in the front-end of the SM process, in the diffusion area of the wafer fab, where dopants are introduced into the silicon wafers at high temperatures to alter their electrical properties. This is achieved using diffusion furnaces that operate in batches, applying constant heat on the wafers for long periods, typically 10-12 hours [17]. These prolonged processing times are essential to ensure the thermal stability required for an even distribution of dopants on the wafers [18]. Due to the chemical nature of the process, not all the jobs can be batched together. Instead, incompatible job families are defined by the distinct “recipes” of the jobs, which dictate unique dopants and thermal conditions [2]. Jobs with the same recipe are said to be compatible and therefore grouped into the same family; jobs in the same family also have the same processing time [1, 19]. Jobs of different families are said to be incompatible and cannot be processed together in the same batch, as this may result in unwanted cross-contamination in the chemical process. Hence, the scheduling problem in the diffusion area belongs to the incompatible case of p-batch, where the batches can only contain jobs of the same family [2, 18]. Furthermore, due to the risk of burning the wafers or damaging the chemical process, jobs cannot remain in the furnace beyond their processing time. Moreover, the processing of a batch cannot be interrupted to include or remove jobs mid-process, as this would disrupt the thermal stability and jeopardize the uniformity of the diffusion [18]. Thus, jobs inside a batch must start and end together.
Existing exact methods in the literature for p-batching include mixed-integer programming (MIP) [6, 18] and Constraint Programming (CP) models [8, 18]. MIP models do not scale well and are outperformed by the CP approaches [2, 18]. However, the existing CP model for p-batching with incompatible jobs is solver-dependent as it relies on specific functions and global constraints only available in a well know commercial CP solver. This paper expands the body of knowledge around this problem by proposing four new CP models that can be implemented in other solvers: one model that relies on automaton constraints, available in different open-source solvers; and other three models integrate assignment and scheduling decisions with different strategies and global constraints, available in many commercial and open-source solvers. Furthermore, the paper conducts thorough computational experiments that demonstrate the implementation flexibility of the proposed models, and compares them under different objectives.
@@
The remainder of this paper is organized as follows. Section II presents a literature review on p-batching. Section III clearly outlines our contributions. Section IV formally presents the description of the problem we address. Section V presents the existing solver-dependent CP model for p-batching. Section VI presents the four proposed CP models. Section VII presents our computational experiments and their results. Finally, Section VIII presents the conclusions and outlines future lines of research.
@@ II Literature Review
The standard classification scheme for scheduling problems provided by Graham et al. [20] is $\alpha|\beta|\gamma$, where $\alpha$ denotes the scheduling environment, $\beta$ describes the job and family characteristics and any restrictions, and $\gamma$ defines the objective function. This section focuses on relevant articles that deal with multiple parallel machines (i.e., $\alpha=P$). For a more comprehensive overview of single-machine p-batch scheduling, interested readers are referred to the recent survey by Fowler and Mönch [2]. However, a notable mention should be made to the CP model by Malapert et al. [21], which solves instances of up to 50 compatible jobs in a single machine in just a few seconds thanks to their implementation of the pack and sequencing global constraints in Choco Solver [22].
@@
For $\beta$, the focus is on articles that deal with multiple incompatible families (i.e., $P|p-batch,incompatible|\gamma$), although key articles with a single compatible family (i.e., $P|p-batch|\gamma$) are also considered, as the paper extends some of their concepts to the incompatible case. The review also includes some articles that consider non-identical job sizes ($s_{j}$) and release times ($r_{j}$). For $\gamma$, the reviewed objective functions primarily fall into three categories: (i) utilization objectives such as minimizing the makespan ($C_{\max}$); (ii) cycle time objectives such as minimizing the total completion time (TCT) and its weighted counterpart (TWCT); and (iii) on-time delivery objectives such as minimizing earliness, lateness, tardiness, number of tardy jobs, including their weighted variants [2]. However, since the diffusion area is usually a bottleneck in the wafer fab, fab leaders often focus on two key aspects: the cycle time of the jobs to increase the effective movement of products [19] and the utilization of the diffusion furnaces. Thus, this paper focuses on minimizing the TWCT, which specifically targets the cycle time performance [2], and the makespan ($C_{\max}$), which specifically targets the utilization performance [2].
@@
Most of the articles that consider multiple parallel machines usually focus on a single family (i.e., $P|p-batch|\gamma$) [2]. This problem is NP-hard [23, 2] and multiple heuristic algorithms [24, 25, 26] as well as meta-heuristics [27, 28, 29] have been proposed. Some exact approaches include mixed-integer linear programming (MILP) models [16] and efficient arc-flow formulations [30, 31, 32, 5]. Recently, CP has been used for this problem in contexts different from SM, such as the composite manufacturing industry [11], and AM processes where parts with irregular shapes are 3D printed [8, 9, 10]. Tang and Beck [11] formulated a CP model and a logic-based benders decomposition (LBBD) for the composites manufacturing industry where the first stage is a compatible serial-batch of layers of raw materials and epoxy resin into mould tools, and the second stage is a compatible p-batch of the mould tools on autoclave ovens. Their CP model cannot find solutions to instances of more than 25 jobs but their LBBD does. In this paper we present four CP models that can find solutions to instances with up to 200 jobs. Nascimento et al. [8, 9] presented an Assign-and-Schedule (AS) CP model that uses binary variables for the assignment of the irregular-shaped parts to printing batches and CP scheduling logic to sequence the batches on the 3D printers. Their assignment logic already locates the parts inside the batch, so they do not explicitly include capacity constraints. However, they only consider compatible jobs (i.e., one family). This paper extends their formulation to handle multiple incompatible families, and include additional constraints to handle the batch size requirements.
@@
The literature that deals with multiple parallel machines and incompatible job families (i.e., $P|p-batch,incompatible|\gamma$) is rather scarce [2]. Since this problem is NP-hard [4], the literature is dominated by heuristics and meta-heuristics [4, 33, 34, 35, 36, 37, 19]. However, exact methods have also been proposed, such as dynamic programming (DP) [19] and MIPs [6, 18]. Cakici et al. [6] proposed a time-indexed (TI) MIP model inspired by the truck delivery problem. They also presented two variable-neighborhood search (VNS) heuristics that outperform the MIP model. Later, Ham et al. [18] presented another MIP formulation based on positional assignment (PA) variables in the context of the diffusion area. They also proposed a CP model that outperforms both the PA and the TI MIP models, as well as the VNS heuristics.
The core idea behind the model by Ham et al. [18] is also presented by Laborie [38][Slide 129] and Laborie et al. [39]. This core idea can be described in three parts. First, the usage of a state function that indicates the family being processed on a machine. This state function is a set of non-overlapping segments over which the function maintains a constant non-negative integer value (i.e., the family being processed). Second, the usage of a specialized alwaysEqual global constraint that ensures that the state of a machine while processing a job is always equal to the job’s family. Third, the usage of two optional alignment parameters in the alwaysEqual constraint. These alignment parameters ensure that if a job is processed on a machine, its start and end times are aligned with the start and end times of the time segment where the machine’s state function takes the constant value. Because of this, the model presented by Ham et al. [18] is henceforward referred to as the Aligned (A) model. This A model performs very well due to the propagation and pruning strategies behind the global constraint and alignment parameters in it. However, the state function, the alwaysEqual constraint, and the two alignment parameters are only present in the commercial solver by IBM, ILOG CP Optimizer [39].
@@ III Contributions
As evidenced in the previous section, most of the existing CP models for p-batch only deal with a single compatible family [8, 9, 10]. The AS model by Cakici et al. [10] relies on a constraint specifically designed to ensure the correct batch duration in the AM context, which does not directly translate to the diffusion problem [10]. Moreover, the A model by Ham et al. [18] relies on a state function, a global constraint, and two extra alignment parameters that are only available in a commercial solver. Thus, the p-batch scheduling literature that considers multiple incompatible job families could benefit from new solver-independent CP models.
This paper addresses these gaps in the literature by proposing four new CP models that consider multiple incompatible job families and can be implemented in different commercial and open-source solvers: (i) A new Automaton (AU) model that relies on automaton constraints, available in different open-source solvers, to handle the progression of the elapsed processing time of a batch. To the best of the authors’ knowledge, this is the first CP model for p-batching that uses this type of constraints in the literature. (ii) An Assign-and-Schedule (AS) model inspired by Nascimento et al. [8, 9] that handles the assignment of jobs to batches with binary variables, and the sequencing of batches on machines using CP scheduling logic. (iii) A Synchronized (S) model inspired by Cakici et al. [10] that replaces the binary assignment with CP scheduling logic, and guarantees the presence of the required batches with an implication constraint. (iv) A Redundant Synchronized (RS) model that improves computational and quality performance through redundant variables and constraints. The thorough computational experiments on standard test cases compare these models using different CP solvers under different objectives and demonstrate the implementation flexibility of the proposed models.
@@ IV Problem Description
Let $\mathcal{J}$ be the set of jobs, $\mathcal{F}$ be the set of incompatible job families, and $\mathcal{M}$ be the set of machines. All jobs can be scheduled on all machines. Compatible jobs belong to the same family. Hence, let $f_{j}\in\mathcal{F}$ be the family of job $j\in\mathcal{J}$, and $\mathcal{J}_{f}=\set{j\in\mathcal{J}:f_{j}=f}$ be the subset of jobs that belong to family $f\in\mathcal{F}$. All the jobs of the same family have the same processing time, thence, let ${\delta}^{f}$ be the processing time of family $f$ and let ${\delta}_{j}={\delta}^{f}$ be the processing time of job $j\in\mathcal{J}_{f}$. Additionally, each job $j$ has a size ${n}_{j}$ (i.e., number of parts), a weight $\omega_{j}$ (i.e., priority), and a release time $r_{j}$. Each machine $m\in\mathcal{M}$ can simultaneously process multiple jobs of the same family $f\in\mathcal{F}$ in batches whose maximum size is $u_{f}$. Jobs in the same batch have to start and end together in a synchronized fashion, and jobs of different families cannot be processed together in the same batch. In the classification by Graham et al. [20], this problem is $P|s_{j},r_{j},p-batch,incompatible|\gamma$. For the objective $\gamma$, the goal is to minimize either the total weighted completion time (TWCT) or the makespan ($C_{\max}$) of the jobs.
V The Existing Aligned Model
@@
Model 1 presents the A model proposed by Ham et al. [18] to schedule incompatible jobs in the diffusion area. Equation (1) defines an interval variable $x_{j}$ of the form $[a,a+{\delta}_{j})$ (i.e., it has a duration of exactly ${\delta}_{j}$ units of time), which represents job $j\in\mathcal{J}$. Note that the start time of this interval must be inside of the scheduling horizon =$\mathcal{H}=\set{0,1,\ldots,H}$, where $H=(\max_{j\in\mathcal{J}}r_{j})+(\sum_{j\in\mathcal{J}}{\delta}_{j})+1$ is the horizon’s limit. Furthermore, this interval can only start after the job’s release time. Equation (1) defines an optional interval variable $x_{jm}$, also of size ${\delta}_{j}$, that represents the option of processing job $j$ on machine $m\in\mathcal{M}$. The presence of the interval $x_{jm}$ is optional since it is allowed to take the value $\perp$, which indicates its absence in the solution. These optional interval variables allow the model to decide which machine should process the jobs. Equation (1) defines a function ${\mathfrak{f}}_{m}$, which is the unique state of the machine $m$ at every point in time. Equation (1) defines a cumulative function ${\mathscr{l}}_{mf}$, which represents the load on machine $m$ of jobs of the family $f\in\mathcal{F}$. It uses the pulse function, which receives two inputs: An interval $x_{jm}$ and an integer value ${n}_{j}$. This pulse function tells the cumulative function to increment its value in ${n}_{j}$ units during the interval $x_{jm}$, if present. Hence, the function ${\mathscr{l}}_{mf}$ accumulates the sizes of overlapping jobs of family $f\in\mathcal{F}$ on machine $m$ at every point in time.
Model 1 Aligned model by Ham et al. [18]
@@
$x_{j}$$\in\{[a,a+{\delta}_{j}):r_{j}\leq a\in\mathcal{H}\},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J};$ $(1)$
$x_{jm}$$\in\{[a,a+{\delta}_{j}):a\in\mathcal{H}\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M};$ $(2)$
@@
${\mathscr{l}}_{mf}$ $=\sum_{j\in\mathcal{J}_{f}}\text{pulse}\left(x_{jm},{n}_{j}\right),$ $\forall\leavevmode\nobreak\ m\in\mathcal{M},f\in\mathcal{F}.$ $(4)$
@@
alwaysEqual $\left({\mathfrak{f}}_{m},x_{jm},f_{j},{\texttt{True},\texttt{True}}\right);$
$\forall\leavevmode\nobreak\ m\in\mathcal{M},f\in\mathcal{F}:\quad{\mathscr{l}}_{mf}\leq u_{f}.$ $(8)$
Objective function (1) minimizes the TWCT. Constraints (1) use the $\texttt{alternative}{(v,V)}$ global constraint, which receives an interval variable $v$, and a set of optional interval variables $V$. This global constraint ensures that, if the interval variable $v$ is present, then only one of the optional interval variables in $V$ gets selected to be present, and $v$ gets synchronized with the selected interval in $V$ [40]. Hence, constraints (1) enforce each job to be processed on exactly one machine. Constraints (1) use the global constraint $\texttt{alwaysEqual}{(h,v,k,align\_s,align\_e)}$, which receives a state function $h$, an interval variable $v$, an integer value ${k}$, and two optional boolean parameters $align\_s$ and $align\_e$. This global constraint ensures that the state function $h=k$ during the interval $v$, if present [40]. Besides, if the two boolean parameters are set to True, the start and end times of the interval $v$ are aligned with the start and end times of the time segment where the state function $h$ takes the value of $k$. Hence, constraints (1) ensure that jobs of only one family are processed at any time on the machines in batches, and that jobs in the same batch start and end together. Finally, constraints (1) impose an upper bound on the cumulative functions that captures the load on the machines by family, ensuring the maximum batch sizes.
To minimize the makespan, objective function (1) should be replaced by objective function (9), which minimizes the maximum end time of the jobs. Although this nonlinear objective function should be linearized in a linear programming (LP) model, a CP solver is able to handle it as presented.
minimize $\text{minimize\leavevmode\nobreak\ \leavevmode\nobreak\ }\max_{j\in\mathcal{J}}\leavevmode\nobreak\ \texttt{endOf}\left(x_{j}\right)$ (9)
Consider the scenario in Table I as an illustrative example with four jobs, one family, and one machine. Assume that the size of each batch must be at most $u_{1}=100$. Figure 1 shows the Gantt chart of different solutions of the A model in different settings. Figure 1(a) shows the solution with these alignment parameters set to True and minimizing the TWCT. Note how the machine state is divided in two segments and the start and end times of the jobs are aligned with the start and end times of their corresponding state segment. Figure 1(b) shows the aligned solution minimizing the makespan. In constrast, Figure 1(c) shows the solution with either objective and the alignment parameters set to False, which allows jobs to start mid-process of other jobs, making this unaligned solution infeasible.
@@
Job $j$ Size ${n}_{j}$ Weight $\omega_{j}$ Release time $r_{j}$ Family $f_{j}$ Processing time ${\delta}_{j}$
@@
(a) TWCT and alignment parameters set to True
(b) $C_{\max}$ and alignment parameters set to True
(c) Alignment parameters set to False
Figure 1: Optimal solutions in different settings
Although the A model is the current state-of-the-art CP model for p-batch scheduling with incompatible job families, the state function (1) and the global constraint (1) are only available in the commercial solver IBM ILOG CP Optimizer, making it solver-dependent. Similar to how an LP model is independent of the solver implementation, the scheduling literature could benefit from new CP models that can be implemented in other commercial and open-source solvers.